10 When the attitude changes from 1.000 to 1.500 the air density changes from. 831 to 678. The is a relationship between attitude and air density 11. Solve for x in the following equation: -18 +4(6-2): 12. Identify the elements of scarcity, choice, and opportunity cost in each of the following The Environmental Protection Agency is considering an order that a 500 acre area on the outskirts of a large city be preserved in its natural state, because the area is home to a rod a housing development on the land. open libray ent that is considered an endangered species. Developers had planned to build (source Principles of Economics University of Minnesot 13. Identify the elements of scarcity, choice, and opportunity cost in each of the following The manager of an automobile assembly plant is considering whether to produce cars or sports utility vehicles (SUVs) next month. Assume that the quantities of labor and other materials required would be the same for either type of production. (source Principles of Economics, University of Minnesota open library) 14. Identify the elements of scarcity, choice, and opportunity cost in each of the following A young man who went to work as a nurses aide after graduating from high school leaves his job to go to college, where he will obtain training as a registered nurse. (source Principles of Economics, University of Minnesota open library)

plz solve these question
Answer #1

10. Inverse

11. X = -18 + 4(6+2)2 = -18 + 4 x 82 = -18 +256 = 238

12. The Land is a scarce resource because it has alternative uses. Here we will have to choose between two alternative uses of land either using as natural habitat for conserving endangered rodents or in building housing development. Here the oppurtunity cost of using Land for preserving rodents is the revenue expected from housing development.

13. Here the factor inputs labour and other materials are scarce because they have alternative uses. We have a choice of using these factors of production in making either cars or SUV. The oppurtunity cost of making cars is the revenue generated from manufacturing SUV and vice versa.